OpenStudy (anonymous):

x^2-6x = 0 solve by factorising pliz?? any help

jimthompson5910 (jim_thompson5910):

factor out x to get x^2 - 6x = x(x-6)

jimthompson5910 (jim_thompson5910):

so x^2-6x = 0 turns into x(x-6) = 0 then use the zero product property to get x = 0 or x-6 = 0

jimthompson5910 (jim_thompson5910):

so x = ??? or x = ???

OpenStudy (anonymous):

Notice that this quadratic can be written as x^2 - 6x + 0 = 0 This implies that we are finding two numbers which multiply to 0 and add up to -6. Obviously, such two numbers are 0 and -6. So we can factor accordingly like this: \[x^2-6x=0 \rightarrow x^2 - 6x + 0 = 0 \rightarrow (x+0)(x-6)\implies x(x-6)=0\] Can you find what x can be from here? @timkainavaka

OpenStudy (anonymous):

ok fanx very much genius and jim for help...now i got it u both are legend....fanx anyway...